Two long parallel wires are 10cm apart and carry currents of 6A and 4A. What is the field B at 4A carrying due to 6A wire.
1
Expert's answer
2020-05-15T08:41:37-0400

The magnetic field created by a long wire with current


"B=\frac{\mu_0 I}{2\pi r}\\
=\frac{4\pi\times 10^{-7}\times 6}{2\pi \times 0.1}=1.2\times 10^{-5} \:\rm T"